OpenStudy (anonymous):

What is the sum or the first five terms of a geometric series with a1=20 amd r=1/4

hartnn (hartnn):

use the sum formula! \(\Large S_n = a_1 \dfrac{r^n -1}{r-1}\) here , a1 = 20 r = 1/4, n= 5 just plug in values :)